Solve the following 1D wave equation: ?tt (?, ?) = ?xx= (?, t) with the boundary conditions ?(0, ?) = ?x(1, ?) = 0, where ?(?, ?) refers to the twist angle of a uniform rod of unit length.
Answer #1

twist angle of Wave Given Olu, t) aufers to the a uniform red rod od unit length O. OAX (2,t) equation : H(24t) - 2?0(2) at²x+x²x=0 X(2) clos (2x) + D sin (on) Bouwselong conditions Olet)-0 a X(6) TH)=0 3 xe)- 0 as T(+) to X(0) = 0 clos (0) to to sand Tuft) = An Los (R-1) set ces (RM-1) ot) + Runsin (22:) st) by primiple of hepezepesition : 8 (m, t) (02:1) st) sin (29.)